Let MC = 8+6Q and MR = 80-10Q. What is the profit maximising output for this firm?
1
Expert's answer
2011-11-18T10:53:33-0500
Let MC = 8+6Q and MR = 80-10Q. What is the profit maximizing output for this firm?
MC=MR
8+6Q=80-10Q 16Q=72& Q=4,5;
Answer:1) Q=2; 2) Q=4,5